OpenStudy (anonymous):

Andy’sweight fluctuates according to a sinusoidal function of time. Let x = 0 correspond to the beginning of the year (January, 1st). Because of his newyear resolutions, Andy loses weight and reaches aminimumof 205 pounds at x = 60 days. Andy then stops exercising and dieting and his weight increases to a maximum of 250 pounds at x = 140 days. If summer starts on the day x = 171 and ends on the day x = 265, how many days of summer will Andy’s weight be below 220 pounds (these are the days he fits into his bathing suit)? So I found the function which is 22.5sin(2pi/160(t-100))+227.5

OpenStudy (anonymous):

So there is a min at x=60 and a max at x=140, how did you go about solving the equation?

OpenStudy (anonymous):

so the form for a sinusoidal function is asin(2pi/b(x-c))+d a=max-min (.5) d= max +min (.5) b= (140-60)(2) c= xmin+b/4 160/4+60

OpenStudy (anonymous):

so type the work out for us to verify

OpenStudy (anonymous):

so the form for a sinusoidal function is asin(2pi/b(x-c))+d a=max-min (.5) d= max +min (.5) b= (140-60)(2) c= xmin+b/4 160/4+60

OpenStudy (anonymous):

a= 250-205/2=22.5 d=250+205/2=227.5 b=140-60(2)=160 c= 60+160/4=100